# Evaluating Impacts of Traffic Regulations in Complex Mobility Systems Using Scenario-Based Simulations

### v1

@@ Abstract
Urban traffic regulation policies are increasingly used to address congestion, emissions, and accessibility in cities, yet their impacts are difficult to assess due to the socio-technical complexity of urban mobility systems. Recent advances in data availability and computational power enable new forms of model-driven, simulation-based decision support for transportation policy design. This paper proposes a novel simulation paradigm for the ex-ante evaluation of both direct impacts (e.g., traffic conditions, modal shift, emissions) and indirect impacts spanning transportation-related effects, social equity, and economic accessibility. The approach integrates a multi-layer urban mobility model combining a physical layer of networks, flows, and emissions with a social layer capturing behavioral responses and adaptation to policy changes. Real-world data are used to instantiate the current “as-is” scenario, while policy alternatives and behavioral assumptions are encoded as model parameters to generate multiple “what-if” scenarios. The framework supports systematic comparison across scenarios by analyzing variations in simulated outcomes induced by policy interventions. The proposed approach is illustrated through a case study aims to assess the impacts of the introduction of broad urban traffic restriction schemes. Results demonstrate the framework’s ability to explore alternative regulatory designs and user responses, supporting informed and anticipatory evaluation of urban traffic policies.
@@ 1 Introduction
The availability of data for sensing and observing urban systems has grown at an unprecedented pace. Advances in pervasive sensing, mobile devices, connected vehicles, and digital platforms now provide access to urban data with increasingly fine-grained spatial, temporal, and social resolution (Zheng, 2019). In parallel, rapid progress in computational capabilities has substantially expanded. Together, these developments are enabling public authorities and urban decision makers to adopt novel, powerful data-driven paradigms, enabling the systematic evaluation of policy interventions beyond what was previously feasible.
@@
At its core, the framework relies on an urban mobility model that explicitly integrates a physical layer, including the transportation network, traffic flows, and pollutant emissions, with a social layer capturing individual and collective behavior, adaptation mechanisms, and varying degrees of behavioral rigidity. This coupling is essential to represent the socio-technical nature of urban mobility systems and to account for behavioral responses to policy interventions.
The “as-is” scenario is instantiated using directly measured and estimated quantities derived from available data sources. These include mobility-related indicators such as vehicle volumes, origin–destination matrices, and emissions, represented at both macro and micro levels. Policy alternatives are then defined through model parameters that encode endogenous events (e.g., public authority regulations), exogenous events (e.g., infrastructure disruptions or changes in vehicle fleets), and assumptions on user responses and behavioral adaptation.
@@
This paper is structured as follows. Section 2 presents the conceptual framework of the study, highlighting the transportation policies and the standard evaluation methods, as well as explaining the challenges in modeling mobility systems. Section 3 outlines the general architecture of the proposed methodological pipeline and its application to a specific case study. Section 4 formalizes the mathematical model constructed for the case study. Section 5 details the implementation and application of the model, reporting the assessed results. Finally, Section 6 summarizes the main findings and discusses the limitations and challenges for future research.
@@ 2.1 Policy application and evaluation
In recent decades, transportation policies have assumed a central role in the international political agenda, driven by the need to balance the traffic externalities with economic growth, social equity and accessibility, and environmental objectives. The literature has identified various types of traffic policy interventions, reflecting different strategic approaches to address transport challenges. Operationally, these policies can be categorized according to whether they target the supply side or the demand side of the transport system.
Supply strategies are designed to increase the quality of the supply and, hence, its traffic capacity and efficiency. Examples of such traffic policies are the modifications to the transport infrastructure, both aimed at equipping the system with new facilities (e.g., the construction of a new highway or a new railway connection), and at enhancing the existing system through maintenance and expansion (e.g., the increase of the capacity of an existing road through the introduction of new lanes). Similarly, the technological enhancement of the supply, towards the constitution of Intelligent Transport Systems, is designed to monitor and manage traffic in real-time and provide information to travelers about traffic conditions, enabling better route choices and reducing congestion. Another key example of supply policy is the optimization of the public transport systems, where the scheduling and the routing are recalibrated to properly balance the mobility demand with the mobility offer. As demonstrated by Rhoads and Shogren (2006), supply-side policies tend to have a higher social discount rate and are often preferred to obtain immediate and tangible changes in the transport system. However, they may induce additional demand and generate the so-called induced traffic, sometimes offsetting long-term benefits (Goodwin, 1996; Downs, 2004; Hymel, 2019).
@@ 3.1 Motivating case: traffic regulation pricing policy
The motivating case considered focuses on an urban area characterized by recurring congestion during peak hours and significant levels of atmospheric emissions. In such contexts, implementing traffic regulation policies is crucial, yet their evaluation proves challenging due to the complex interactions between user behavior, network characteristics, and regulatory parameters. The policy under study is a traffic pricing system, where tolls vary according to temporal and categorical criteria, with exemptions for specific user groups and a defined application schedule.
@@ 3.2 General framework architecture
Figure 1: Schematic representation of the architectural framework for what-if scenario analysis
@@
In this framework, variables can be defined either deterministically, with fixed values, or stochastically, as random variables following probability distributions. The probabilistic formulation allows us to explicitly represent input uncertainty and trace how it propagates through the model to the final estimates of the quantities of interest. By treating parameters probabilistically rather than as fixed values, the model can be evaluated repeatedly, with each run producing potentially different outputs that reflect the underlying uncertainty. An ensemble approach enables the quantification of output uncertainty, either through small variations in input values or, in the case of nondeterministic models, by using different seeds for the random processes.
@@ 3.3 Application to the case study
The policy enacts a area-based traffic regulation for which the vehicular circulation in a designated area is limited. In this example, the regulation imposes a fee on all vehicles circulating into the area during the daily hours. The cost depends on the type of vehicle in circulation: to achieve emission reduction goals alongside traffic reduction, different penalty levels can be applied to vehicles based on their emission standards, for instance, with higher costs assigned to older, more polluting classes. Moreover, the policy provides payment exemptions for vulnerable social groups.
@@
The application of the policy and its effects on the user travel behaviours initiate a sequential process that progressively alters inflow, traffic, and emission levels in the modified scenario. The effects on the inflow follow directly from the policy implementation. In the new scenario, the inflow varies from the base one with time: it may be reduced when the policy is applied, and there remain only exempted and rigid travelers; on the other hand, it may be increased in the hours of policy inactivation, when anticipating and postponing travelers are to be added to the base inflow. When aggregating traffic from all vehicles, both those starting their journey inside the area and those arriving from outside, the total traffic volume changes in accordance with the variations in the number of vehicles. The new scenario is hence associated with a modified value of total traffic in the area per time instant. This causes changes in the vehicular emissions from the area, which vary according to the traffic. The policy and the subsequent scenario modification affect other measurable quantities which could be important for a complete policy assessment. For instance, within the transport domain, the extent of the modal shift towards public transport can be assessed.
@@ 4.1 Scenario definition
Among the quantities of interest, $I(t)$ represents the vector of inflows, that is, the number of vehicles entering the area in each time interval $t$. In the current base scenario, these inflow values are treated as known parameters, obtained through estimation procedures applied to real traffic data, while in the modified scenario, their values are estimated with the application of the behavioral models explained in Section 4.2. The other two core variables are $T(t)$, representing the number of circulating vehicles in the area during time interval $t$, and $E(t)$, which denotes the emissions produced by all vehicles circulating within the area during time interval $t$. The estimation of these quantities is presented in Section 4.3 and Section 4.4, respectively, and follows from the known inflow and starting patterns combined with the emission level distribution of the active vehicle fleet.
@@ 4.2 Inflow modifications
Similarly, Equation 5 uses the acceptance anticipation $\Delta_{a}$, the required anticipation per time interval $\Delta_{fs}(t)$, and the median parameter $\Delta_{a50\%}$, to compute the probability of accepting anticipation. However, to fully avoid the circulation policy, the trip should be anticipated to be completed before the policy start. Consider the travel time in the area, modeled as a random variable with a geometric distribution $\tau\sim\mathcal{G}eom(p=\tfrac{1}{\bar{\tau}})$ over the support $\mathbb{N}_{>0}$. Then the probability of anticipation can be found as in Equation 5.
$\mathbb{P}[\text{anticipation in time }t]=\sum_{T=1}^{\infty}\mathbb{P}[\Delta_{a}\geq\Delta_{fs}(t)+\tau|\tau=T]\cdot\mathbb{P}[\tau=T]=\varepsilon\cdot\exp\left\{-\frac{\ln(2)}{\Delta_{a50\%}}\cdot\Delta_{fs}(t)\right\}$ (5)
where $\varepsilon=p\,e^{-{\ln 2}/{\Delta_{a}50\%}}[1-(1-p)e^{-{\ln 2}/{\Delta_{a}50\%}}]^{-1}$.
@@
Two additional indices, $x_{z3}$ and $x_{z4}$, evaluate the convenience of public transport compared to private transport: cost and time difference. The cost of the public transportation service is the fare that a user would pay to travel via public transportation from any zone to the area of interest, evaluating the economic convenience of the mode-shift. The difference in travel times when using public transport and when using private means, then, assess the temporal convenience of the mode-shift.
@@
By reweighting the probability with the proportion of vehicles starting in each zone, $Z_{z}$, the probability of overall modal shift acceptance is found, as in Equation 8.
@@
This model accounts a time-shifting strategy that separates the shifting decision into two stages. First, travelers decide whether to accept time-shifting based solely on avoiding the policy period, regardless of the specific alternative timing. Then, those who accept are redistributed to new departure times before or after the policy period without considering their original departure preferences.
@@
The absolute number of anticipated redistributed vehicles at each time interval is then computed as the fraction $F_{ar}(t)$ multiplied by the total anticipating $TN_{a}$. Similarly, the number of postponed redistributed vehicles is assessed multiplying $F_{pr}(t)$ by $TN_{p}$.
@@
Given the proportion of vehicles implementing each behavioral strategy, it is possible to calculate the number of vehicles entering the policy-affected area under the modified scenario. Equation 14 presents the formulation for computing the inflow, which distinguishes between two cases depending on policy application time. Within the period of enaction of the policy, with $t\in[t_{s},t_{e}]$, the inflow is the sum of exempted vehicles and rigid vehicles. Outside the hours of the policy, with $t\notin[t_{s},t_{e}]$, the inflow is modified from a portion of vehicles that anticipate and postpone the starting time and sum to the original vehicles.
@@ 4.3 Traffic estimation model
The objective is to compute the traffic vector satisfying Equation 15. Although this system admits a direct solution, we adopt an iterative approach as outlined in Algorithm 1. The direct and iterative approaches are mathematically equivalent and yield identical solutions; however, the iterative formulation empirically assesses superior computational performance, particularly in terms of execution time, and is hence chosen for the implementation . The starting point is an initial guess for the traffic values, $T(t)^{0}=I(t)+S(t)$ for all $t$. Then, at each iteration step $k$, the algorithm recomputes the value of the retention probability $\alpha^{k}$ and updates the traffic estimate $T^{k+1}$, as shown in Equation 16.
@@
Having described the traffic model, we now address the computation of its input data, which are required to estimate the traffic vectors in both the reference and policy scenarios. First, in the reference scenario, $T$ is computed using the inflow and starting vectors $I$ and $S$. The inflows $I$ and the starting trips $S$ are extracted from observed data representing the current situation. Then, in the policy scenario, the modified traffic $T_{m}$ is estimated starting from the corresponding modified vectors $I_{m}$ and $S_{m}$, where $I_{m}$ is adjusted according to the procedures outlined in the previous subsection. For the starting vector, the same reasoning applied to the inflow can be adopted. All vehicles that start within the area of interest, excluding exempted vehicles, are required to pay a ticket fee and are therefore directly affected by the policy. As a result, they may decide whether and how to adjust their mobility habits with rigidity, time-shifting, mode-shifting, or cancelation. The modified starting vector was computed following the same procedure used for the inflow case, replacing the inflow vector $I$ with the starting vector $S$.
@@ 4.4 Emission estimation model
For the calculation of emissions in the area, it is assumed that each vehicle in circulation produces a value of average emissions per kilometer depending on its characteristics. This work uses the classification of vehicles in Euro emission standards (Selleri et al., 2021), widely adopted in European environmental and transport studies, to standardize the assessment of vehicle emissions and ensure comparability across different regions and datasets. Vehicles are classified into eight Euro classes, from Euro-0 to Euro-$7$, based on their characteristics such as vehicle type, fuel type, date of the first registration, and maximum emissions per kilometer. Accordingly, we define $\bar{E}_{l,km}$ as the representative average emissions per kilometer corresponding to the $l$-th Euro-class, with values obtained from the literature. Moreover, let $P_{l}$ represent the percentage of vehicles for each of the Euro-class $l$. This proportion reflects the actual composition of the fleet in the current base scenario, derived from the elaboration of real-world data.
@@
Then, it is possible to compute the Euro-level distribution under the policy enactment. During the period $t_{s},t_{e}]$, the new proportion $P_{l,m}(t)$ accounts for two contributions. First, it considers the fraction of exempt vehicles $P_{l}\cdot F_{e}$, which represents the part of the Euro-class $l$ that is not subject to the policy and remains unchanged. Second, it consider the fraction of rigid vehicles, $P_{l}\cdot\frac{F_{rigidity,l}(t)}{F_{e}+F_{rigidity}(t)}$, which represents the proportional contribution of the rigid share of class $l$ compared to the total of rigid vehicles and exempt vehicles. Equation 18 formalizes this reasoning.

### v2

@@ Abstract
Urban traffic regulation policies are increasingly used to address congestion, emissions, and accessibility in cities, yet their impacts are difficult to assess due to the socio-technical complexity of urban mobility systems. Recent advances in data availability and computational power enable new forms of model-driven, simulation-based decision support for transportation policy design. This paper proposes a novel simulation paradigm for the ex-ante evaluation of both direct impacts (e.g., traffic conditions, modal shift, emissions) and indirect impacts spanning transportation-related effects and economic accessibility. The approach integrates a multi-layer urban mobility model combining a physical layer of mobility flows and emissions with a social layer capturing behavioral responses and adaptation to policy changes. Real-world data are used to instantiate the current “as-is” scenario, while policy alternatives and behavioral assumptions are encoded as model parameters to generate multiple “what-if” scenarios. The framework supports systematic comparison across scenarios by analyzing variations in simulated outcomes induced by policy interventions. The proposed approach is illustrated through a case study that aims to assess the impacts of the introduction of broad urban traffic restriction schemes. Results demonstrate the framework’s ability to explore alternative regulatory designs and user responses, supporting informed and anticipatory evaluation of urban traffic policies.
@@ 1 Introduction
The availability of data for sensing and observing urban systems has grown at an unprecedented pace. Advances in pervasive sensing, mobile devices, connected vehicles, and digital platforms now provide access to urban data with increasingly fine-grained spatial, temporal, and social resolution (Zheng, 2019). In parallel, rapid progress in computational capabilities has substantially expanded. Together, these developments are enabling public authorities and urban decision makers to adopt novel, powerful data-driven paradigms, allowing the systematic evaluation of policy interventions beyond what was previously feasible.
@@
At its core, the framework relies on an urban mobility model that explicitly integrates a physical layer, including the mobility traffic flows and pollutant emissions, with a social layer capturing individual and collective behavior, adaptation mechanisms, and varying degrees of behavioral rigidity. This coupling is essential to represent the socio-technical nature of urban mobility systems and to account for behavioral responses to policy interventions.
The “as-is” scenario is instantiated using directly measured and estimated quantities derived from available data sources. These include mobility-related indicators such as vehicle volumes, origin–destination matrices, and emissions, represented at both macro and micro levels. Policy alternatives are then defined through model parameters that encode endogenous events (e.g., public authority regulations), exogenous events (e.g., infrastructure disruptions, changes in vehicle fleets), and assumptions on user responses and behavioral adaptation.
@@
This paper is structured as follows. Section 2 presents the conceptual framework of the study, highlighting the transportation policies and the standard evaluation methods, as well as explaining the challenges in modeling mobility systems. Section 3 outlines the general architecture of the proposed methodological pipeline and its application to the specific case study of the city of Bologna. Section 4 formalizes the mathematical model constructed for the case study. Section 5 details the implementation and application of the model, reporting the assessed results. Finally, Section 6 summarizes the main findings and discusses the limitations and challenges for future research.
@@ 2.1 Policy application and evaluation
In recent decades, transportation policies have assumed a central role in the international political agenda, driven by the need to balance the traffic externalities with economic growth, social equity, accessibility, and environmental objectives. The literature has identified various types of traffic policy interventions, reflecting different strategic approaches to address transport challenges. Operationally, these policies can be categorized according to whether they target the supply side or the demand side of the transport system.
Supply strategies are designed to increase the quality of the supply and, hence, its traffic capacity and efficiency. Examples of such traffic policies are the modifications to the transport infrastructure, both aimed at equipping the system with new facilities (e.g., the construction of a new highway or a new railway connection) and at enhancing the existing system through maintenance and expansion (e.g., the increase of the capacity of an existing road through the introduction of new lanes). Similarly, the technological enhancement of the supply, towards the constitution of Intelligent Transport Systems, is designed to monitor and manage traffic in real-time and provide information to travelers about traffic conditions, enabling better route choices and reducing congestion. Another key example of supply policy is the optimization of the public transport systems, where the scheduling and the routing are recalibrated to properly balance the mobility demand with the mobility offer. As demonstrated by Rhoads and Shogren (2006), supply-side policies tend to have a higher social discount rate and are often preferred to obtain immediate and tangible changes in the transport system. However, they may induce additional demand and generate the so-called induced traffic, sometimes offsetting long-term benefits (Goodwin, 1996; Downs, 2004; Hymel, 2019).
@@ 3.1 Motivating case: traffic regulation pricing policy
The motivating case considers an urban area characterized by recurring congestion during peak hours and significant levels of atmospheric emissions. In such contexts, implementing traffic regulation policies is crucial, yet their evaluation proves challenging due to the complex interactions between user behavior, network characteristics, and regulatory parameters. The policy under study is a traffic pricing system, where tolls vary according to temporal and categorical criteria, with exemptions for specific user groups and a defined application schedule.
@@ 3.2 General framework architecture
Figure 1: Schematic representation of the architectural framework for “what-if” scenario analysis
@@
In this framework, variables can be defined either deterministically, with fixed values, or stochastically, as random variables following probability distributions. The probabilistic formulation allows us to explicitly represent input uncertainty and trace how it propagates through the model to the final estimates of the quantities of interest. By treating parameters probabilistically rather than as fixed values, the model can be evaluated repeatedly, with each simulation producing potentially different outputs that reflect the underlying uncertainty. An ensemble approach enables the quantification of output uncertainty, either through small variations in input values or, in the case of nondeterministic models, by using different initializations for the random processes.
@@ 3.3 Application to the case study
The policy enacts an area-based traffic regulation for which the vehicular circulation in a designated area is limited. In this example, the regulation imposes a fee on all vehicles circulating into the area during the daily hours. The cost depends on the type of vehicle in circulation: to achieve emission reduction goals alongside traffic reduction, different penalty levels can be applied to vehicles based on their emission standards, for instance, with higher costs assigned to older, more polluting classes. Moreover, the policy provides payment exemptions for vulnerable social groups.
@@
The application of the policy and its effects on the user travel behaviours initiate a sequential process that progressively alters inflow, traffic, and emission levels in the modified scenario. The effects on the inflow follow directly from the policy implementation. In the new scenario, the inflow varies from the base one with time: it may be reduced when the policy is applied, and there remain only exempted and rigid travelers; on the other hand, it may be increased in the hours of policy inactivation, when anticipating and postponing travelers are to be added to the base inflow. When aggregating traffic from all vehicles, both those starting their journey inside the area and those arriving from outside, the total traffic volume changes in accordance with the variations in the number of vehicles. The new scenario is hence associated with a modified value of total traffic in the area per time instant. This causes changes in the vehicular emissions from the area, which vary according to the traffic. The policy and the subsequent scenario modification affect other measurable quantities, which could be important for a complete policy assessment. For instance, within the transport domain, the extent of the modal shift towards public transport can be assessed.
@@ 4.1 Scenario definition
Among the quantities of interest, $I(t)$ represents the vector of inflows, that is, the number of vehicles entering the area in each time interval $t$. In the current base scenario, these inflow values are treated as known parameters, obtained through estimation procedures applied to real traffic data, while in the modified scenario, their values are estimated with the application of the behavioral models explained in Section 4.2. The other two core variables are $T(t)$, representing the number of circulating vehicles in the area during time interval $t$, and $E(t)$, which denotes the emissions produced by all vehicles circulating within the area during time interval $t$. The estimation of these quantities is presented in Section 4.3 and Section 4.4, respectively, and follows from the known inflow patterns combined with the emission level distribution of the active vehicle fleet.
@@ 4.2 Inflow modifications
Similarly, Equation 5 uses the acceptance anticipation $\Delta_{a}$, the required anticipation per time interval $\Delta_{fs}(t)$, and the median parameter $\Delta_{a50\%}$, to compute the probability of accepting anticipation. However, to fully avoid the circulation policy, the trip should be anticipated to be completed before the policy starts. Consider the travel time in the area, modeled as a random variable with a geometric distribution $\tau\sim\mathcal{G}eom(p=\tfrac{1}{\bar{\tau}})$ over the support $\mathbb{N}_{>0}$. Then the probability of anticipation can be found as in Equation 5.
[…]
where $\varepsilon=p\,e^{-{\ln(2)}/{\Delta_{a50\%}}}[1-(1-p)e^{-{\ln(2)}/{\Delta_{a50\%}}}]^{-1}$.
@@
Two additional indices, $x_{z3}$ and $x_{z4}$, evaluate the convenience of public transport compared to private transport: cost and time difference. The cost of the public transportation service is the fare that a user would pay to travel via public transportation from any zone to the area of interest, evaluating the economic convenience of the mode-shift. The difference in travel times when using public transport and when using private means, then, assesses the temporal convenience of the mode-shift.
@@
By reweighting the probability with the proportion of vehicles starting in each zone, $Z_{z}$, the probability of overall modal shift acceptance is found as in Equation 8.
@@
This model accounts for a time-shifting strategy that separates the shifting decision into two stages. First, travelers decide whether to accept time-shifting based solely on avoiding the policy period, regardless of the specific alternative timing. Then, those who accept are redistributed to new departure times before or after the policy period without considering their original departure preferences.
@@
The absolute number of anticipated redistributed vehicles at each time interval is then computed as the fraction $F_{ar}(t)$ multiplied by the total anticipating $TN_{a}$. Similarly, the number of postponed redistributed vehicles is assessed by multiplying $F_{pr}(t)$ by $TN_{p}$.
@@
Given the proportion of vehicles implementing each behavioral strategy, it is possible to calculate the number of vehicles entering the policy-affected area under the modified scenario. Equation 14 presents the formulation for computing the inflow, which distinguishes between two cases depending on policy application time. Within the period of enactment of the policy, with $t\in[t_{s},t_{e}]$, the inflow is the sum of exempted vehicles and rigid vehicles. Outside the hours of the policy, with $t\notin[t_{s},t_{e}]$, the inflow is modified from a portion of vehicles that anticipate and postpone the starting time and sum to the original vehicles.
@@ 4.3 Traffic estimation model
The objective is to compute the traffic vector satisfying Equation 15. Although this system admits a direct solution, we adopt an iterative approach as outlined in Algorithm 1. The direct and iterative approaches are mathematically equivalent and yield identical solutions; however, the iterative formulation empirically assesses superior computational performance, particularly in terms of execution time, and is hence chosen for the implementation. . The starting point is an initial guess for the traffic values, $T(t)^{0}=I(t)+S(t)$ for all $t$. Then, at each iteration step $k$, the algorithm recomputes the value of the retention probability $\alpha^{k}$ and updates the traffic estimate $T^{k+1}$, as shown in Equation 16.
@@
Having described the traffic model, we now address the computation of its input data, which are required to estimate the traffic vectors in both the reference and policy scenarios. First, in the reference scenario, $T$ is computed using the inflow and starting vectors $I$ and $S$. The inflows $I$ and the starting trips $S$ are extracted from observed data representing the current situation. Then, in the policy scenario, the modified traffic $T_{m}$ is estimated starting from the corresponding modified vectors $I_{m}$ and $S_{m}$, where $I_{m}$ is adjusted according to the procedures outlined in the previous subsection. For the starting vector, the same reasoning applied to the inflow can be adopted. All vehicles that start within the area of interest, excluding exempted vehicles, are required to pay a ticket fee and are therefore directly affected by the policy. As a result, they may decide whether and how to adjust their mobility habits with rigidity, time-shifting, mode-shifting, or cancellation. The modified starting vector was computed following the same procedure used for the inflow case, replacing the inflow vector $I$ with the starting vector $S$.
@@ 4.4 Emission estimation model
For the calculation of emissions in the area, it is assumed that each vehicle in circulation produces a value of average emissions per kilometer depending on its characteristics. This work uses the classification of vehicles in Euro emission standards (Selleri et al., 2021), widely adopted in European environmental and transport studies, to standardize the assessment of vehicle emissions and ensure comparability across different regions and datasets. Vehicles are classified into eight Euro classes, from Euro-0 to Euro-6, based on their characteristics such as vehicle type, fuel type, date of the first registration, and maximum emissions per kilometer. Accordingly, we define $\bar{E}_{l,km}$ as the representative average emissions per kilometer corresponding to the $l$-th Euro-class, with values obtained from the literature. Moreover, let $P_{l}$ represent the percentage of vehicles for each of the Euro-class $l$. This proportion reflects the actual composition of the fleet in the current base scenario, derived from the elaboration of real-world data.
@@
Then, it is possible to compute the Euro-level distribution under the policy enactment. During the period $t_{s},t_{e}]$, the new proportion $P_{l,m}(t)$ accounts for two contributions. First, it considers the fraction of exempt vehicles $P_{l}\cdot F_{e}$, which represents the part of the Euro-class $l$ that is not subject to the policy and remains unchanged. Second, it considers the fraction of rigid vehicles, $P_{l}\cdot\frac{F_{rigidity,l}(t)}{F_{e}+F_{rigidity}(t)}$, which represents the proportional contribution of the rigid share of class $l$ compared to the total of rigid vehicles and exempt vehicles. Equation 18 formalizes this reasoning.